Allison:

6(4)^2x+8 is rewritten as 6(r)^x+4. What is the value of r?

Allison:

@Hero

Hero:

\(6(4)^{2x+8} = 6(r)^{x+4}\) ?

Allison:

Yep

Hero:

Well, what happens if you factor out 2 from 2x + 8?

Hero:

Even clearer if you divide out the 6 first on both sides.

Allison:

x+8?

Allison:

So 6 is gone from both sides

Hero:

If you divide out the 6 on both sides you end up with \( 4^{2x+8}=r^{x+4}\)

Allison:

Yes

Hero:

So from there you would factor out the 2 from 2x + 8

Allison:

How

Hero:

Question: What happens if you apply the distributive rule to 3(x + 2) ?

Allison:

Rainbow multiply?

Hero:

Never heard of that, but what result do you get if you apply the distributive rule?

Allison:

3x+6?

Hero:

Correct, now what happens if you reverse that process you just performed?

Allison:

Reverse how

Hero:

By factoring.

Allison:

Omg how do I factor

Hero:

I gave you 3(x + 2). You performed the distributive rule to get 3x + 6.

Allison:

Right, rainbow method

Hero:

You got to that result by multiplying the x and the 2 by three.

Hero:

What is the inverse of multiplication?

Allison:

Division

Hero:

exactly

Allison:

What do I divide

Hero:

So when you factor, you're performing division.

Allison:

Right

Hero:

You divide 3x + 6 by 3 and then you write the result of that, then put the divisor on the outside of the parentheses.

Allison:

It's just 6

Hero:

So when you divide 3x+6 by 3, the result is x + 2. And then you put the three on the outside of the parentheses.

Allison:

!

Hero:

So you write 3(x + 2) which is what we started with.

Allison:

Ok

Allison:

What was the point of just going back to where we started?

Hero:

To show you how to factor

Allison:

Ok

Hero:

You asked "how do you factor"

Allison:

Ok now the actual doing of the problem

Hero:

That requires factoring 2x + 8

Hero:

Which requires knowing how to factor

Allison:

2(x+4)

Hero:

Congratulations. You did it.

Allison:

Thanks to you. Now what?

Hero:

I just hope you did it on your own strength.

Hero:

Now you have: \(4^{2(x+4)}=r^{x+4} \)

Allison:

You don't need strength to divide

Allison:

Yes

Hero:

Now it looks like we might be able to say what r is

Allison:

16?

Hero:

Correct

Allison:

Wow how easy